[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Apr 05, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Apr 5 16:18:23 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:

HSU:

Other items:
The GFS continues to load slowly due to the recent network modification. We  
will probably have to live with the slowness until NCO completes the  
process to roughly double our bandwidth by the end of May. Please continue  
to report any new issues through the shift logs to provide us with  
documentation so we can keep the pressure on through the May  
implementation. Also, NCO said they would look at the current network  
settings before the upgrade to see if there are any modifications that  
might help us in the short term.

Our AWIPS2 system has not been able to keep up with the ingest of GOES-16  
data. We may have to turn that off or reduce the number of bands ingested  
until we can find a solution. All the GOES-16 channels will continue to be  
available in N-AWIPS.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170405/afb2a1bf/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list